Find the equilibrium price and equilibrium quantity from the following relations describes demand and supply conditions in a given industry.
QD = 80000 –20000P (Demand)
QS = -20000 + 20000P (Supply)
Where Q is quantity and P is price in dollars.
1
Expert's answer
2020-06-09T17:04:52-0400

At Equilibrium supply is equal demand,


To work work out Price and Quantity at Equilibrium we equate supply to demand as follows;


"QD=80,000-20,000P=QS=-20,000+20,000P"


"80,000-20,000P=-20,000+20,000P"


Equate the likes together,

Price at Equilibrium,PE


"80,000+20,000 =20,000P+20,000P"

"100,000=40,000P"

"PE=2.50\/"


To calculate Quantity we can use one of the equations since we have the value of price

Quantity at Equilibrium,QE


"QD=80,000- ( 20,000*2.5)"

"QE =30,000"